OpenStudy (anonymous):

what does sin^2t =? In relation to trigonometric functions.

OpenStudy (anonymous):

Supposedly it equals 1-cos2t/2 but I don't understand why.

OpenStudy (anonymous):

oh it's just a trig identity i think.

OpenStudy (superdavesuper):

yes it is...do u know cos(A+B)=cosAcosB-sinAsinB? u will need to use that...

OpenStudy (anonymous):

yeah Sin(A+B) = SinACosB + CosASinB

OpenStudy (superdavesuper):

so cos2t=cos(t+t)=....?

OpenStudy (yttrium):

You will use here the double angle identity. Wherein cos2t = 1-2sin^2 t Hence, 2sin^2 = 1 - cos^2 t hence sin^2 t = (1 - cos^2 t)/2

OpenStudy (anonymous):

super it would be cosTcosY - sinTsinY